# Quasi-Monte Carlo with one categorical variable

### v1

@@ Abstract
We study randomized quasi-Monte Carlo (RQMC) estimation of a multivariate integral where one of the variables takes only a finite number of values. This problem arises when the variable of integration is drawn from a mixture distribution as is common in importance sampling and also arises in some recent work on transport maps. We find that when integration error decreases at an RQMC rate that it is then beneficial to oversample the smallest mixture components instead of using a proportional allocation. We also find that for the most accurate RQMC sampling methods, it is advantageous to arrange that our $n=2^{m}$ randomized Sobol’ points split into subsample sizes that are also powers of $2$.
@@ 1 Introduction
We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values. The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios. That is $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$. The distributions $P_{\ell}$ are commonly called mixture components. We call them ‘strata’ to avoid confusion with vector components.
@@
An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$. That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated. Such negative correlation need not happen; we discuss them briefly at the end of the paper.
For some problems the integrand is a simple function within each stratum. A more interesting expression arises in mixture importance sampling. There we estimate $\mu=\int g(\bm{x})p(\bm{x})\mathrm{\,d}\bm{x}$ for a probability density function $p$ by
@@
RQMC does best under favorable regularity assumptions, such as continuity, on the integrand. The presence of the categorical variable makes the resulting integrand discontinuous. That discontinuity is however axis parallel and so it is a QMC-friendly discontinuity in the terminology of Wang and Tan (2013). In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$. Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016). This rate can be attained at powers of $2$ by scrambled Sobol’ points. Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate. By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$. This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within stratum $\ell$ and hence overall.
This paper is organized as follows. Section 2 introduces our notation, gives some background on RQMC, and proves some results about within-stratum discrepancies. Section 3 describes how to allocate samples to strata. In RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes. Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$. Section 5 establishes some minimax variance results from using equal or nearly equal sample sizes for the strata. Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best. Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sample sizes do best. Section 8 has some discussion and conclusions. The proof of one of our results is in Appendix A.
@@
Klebanov and Sullivan (2023) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$. They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path. For $\ell<L$ they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$. Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$. This is different from our strategy of using the first RQMC component to select the stratum which we prefer because it avoids giving two within-stratum samplers the same points. They do not require $n\gg L$. Some of our expressions only require $n>2L$, but as we use asymptotic variances within strata, our results are most relevant for large $n/L$. They consider more general location-scale families than the Gaussian distribution and they can even include some $\alpha_{\ell}<0$ so long as $P$ never has a negative probability density function.
@@ 2.1 Integration variables
We will use RQMC points $\bm{z}\in[0,1]^{s+1}$. We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}$ from $P_{\ell}$. That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$. For a function $\bm{\ell}:[0,1]\to\{1,2,\dots,L\}$, we will sample the $i$’th point using sample stratum $\bm{\ell}(v_{i})$. The pre-image of each $\ell\in\{1,2,\dots,L\}$ under $\bm{\ell}$ will always be a sub-interval of $[0,1]$. We let $\beta_{\ell}$ be the length of that subinterval. We abbreviate $\bm{\ell}(v_{i})$ to $\ell(i)$ in some expressions. Having selected stratum $\ell(i)$, we draw $\bm{x}_{i}\in^{d}$ via $\bm{x}_{i}=\phi_{\ell(i)}(\bm{u}_{i})$.
@@
is a within-stratum sample mean. For the RQMC method we consider we show below that $n_{\ell}\geqslant 1$ whenever $\beta_{\ell}>2/n$. We do not need $\Pr(n_{\ell}>0)=1$ for unbiasedness of $\hat{\mu}$, but without it the stratum sample means are not always well defined.
At this point we can introduce a small generalization. Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
@@ 2.2 Sampling the strata
We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’. By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$. The next proposition shows that if every $\beta_{\ell}>2/n$ then every $n_{\ell}>0$ when $v_{i}$ are stratified.
@@
Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$. Then $\beta=t/n+c$ for $t=r-s$ and $-1/n<c<1/n$. If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
@@ 2.4 Digital nets
For integers $m\geqslant t\geqslant 0$ the $n=2^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$. Smaller $t$ are better. We always have $t\leqslant m$ because $t=m$ just means that all the points are in $[0,1)^{d}$.
@@ 2.5 Discrepancy within each stratum
Here we consider how uniformly distributed the points in each stratum are. The first case we consider has $\beta_{\ell}$ decreasing with $\ell$ and equal to negative powers of $2$. We show how to find such $\beta_{\ell}$ in Section 4. Then the sample points within any given stratum are scrambled nets, and the error can be as good as $O(n^{-3+\epsilon})$ for regular integrands. After that, we consider arbitrary $\beta_{\ell}>2/n$. Then the points in a given stratum still have small discrepancy but they are not generally scrambled digital nets. In that case the rate in stratum $\ell$ is $O(n_{\ell}^{-2+\epsilon})$ for within-stratum integrands that are in BVHK.
@@
If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$. That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$
@@
If $\bm{z}_{1},\dots,\bm{z}_{n}$ are Halton points (scrambled or otherwise) with $n=2^{m}$ then $\bm{z}_{i1}$ for $1\leqslant i\leqslant n$ are stratified.
@@ 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the covergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases. It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer. That can happen when the practically usable values of $n$ are smaller than the sample size at which the asymptotic result holds. Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$. Those often appear to have $21<\rho<$ 2.
@@
With $n_{\ell}=n\beta_{\ell}$ and $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (5) as
@@
The importance sampling ratio $\omega_{\ell}$ has canceled from the expression for $\hat{\mu}$, leaving a simply weighted some of within-stratum means.
@@ 3.1 Optimal sample sizes
RQMC points have a space filling property. We expect that the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$. We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$. This motivates a different criterion
@@
Higher convergence rates $\rho$ tend to equalize the $n_{\ell}$. We can think of $n_{\ell}=n/L$ as arising in an $\rho\to\infty$ limit. We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzs 00 and $1$, and three different hypotheses on $\rho$. They lead to the four different rates given in Table 1. The case for ansatz 00 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation. The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$. This is even true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) don’t distinguish betweem different values of $\rho$. After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
When we do have information about the relative sizes of $\tau_{\ell}$, then we can replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}$ in the formulas in Table 1. In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934). When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$ then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}(\tau_{\ell}/c_{\ell})^{1/2}$ in Table 1.
@@
If $\rho$ varies from stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$. Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.

### v2

@@ Abstract
We study randomized quasi-Monte Carlo (RQMC) estimation of a multivariate integral where one of the variables takes only a finite number of values. This problem arises when the variable of integration is drawn from a mixture distribution as is common in importance sampling and also arises in some recent work on transport maps. We find that when integration error decreases at an RQMC rate that it is then beneficial to oversample the smallest mixture components instead of using a proportional allocation. The optimal allocations depend on the possibly unknown convergence rates. Designing the sample with an incorrect assumption on the rate still attains that convergence rate, with an inferior implied constant. The penalty for using a pessimistic rate is typically higher than for using an optimistic one. We also find that for the most accurate RQMC sampling methods, it is advantageous to arrange that our $n=2^{m}$ randomized Sobol’ points split into subsample sizes that are also powers of $2$.
@@ 1 Introduction
We study randomized quasi-Monte Carlo (RQMC) integration in a setting where one of the input variables is categorical, taking only some finite number $L\geqslant 2$ of different values. The main motivation for this problem is mixture sampling where the variables of interest come from $L$ different distributions representing $L$ different scenarios. That is, $\bm{x}\sim P_{\ell}$ with probability $\alpha_{\ell}>0$ where $\sum_{\ell=1}^{L}\alpha_{\ell}=1$. The distributions $P_{\ell}$ are commonly called mixture components. We call them ‘strata’ to avoid confusion with vector components. We want to estimate $\mu={\mathbb{E}}(g(\bm{x}))=\sum_{\ell=1}^{L}\alpha_{\ell}{\mathbb{E}}(g(\bm{x})\!\mid\!\bm{x}\sim P_{\ell})$. Our approach will also allow us to use distinct functions $g_{\ell}$ within each stratum.
@@
An alternative to the standard MC workflow is to compute $L$ different independently sampled RQMC estimates, one per stratum, and combine them according to their mixture probabilities $\alpha_{\ell}$. That approach misses a potential variance reduction from RQMC sampling when the within-stratum estimates are negatively correlated. Such negative correlation need not happen; we give a reason to expect that at the end of the paper.
For some problems, the integrand is a simple function within each stratum. A more interesting expression arises in mixture importance sampling. There we estimate $\mu=\int g(\bm{x})p(\bm{x})\mathrm{\,d}\bm{x}$ for a probability density function $p$ by
@@
RQMC does best under favorable regularity assumptions, such as continuity, on the integrand. The presence of the categorical variable makes the resulting integrand discontinuous. That discontinuity is however axis parallel and so it is a QMC-friendly discontinuity in the terminology of Wang and Tan (2013). In some very regular cases, RQMC can attain a mean squared error (MSE) of $O(n^{-3+\epsilon})$. Such a rate can only be attained over a sequence of special sample sizes that grow at least exponentially (Sobol’, 1998; Owen, 2016). This rate can be attained by scrambled Sobol’ points with sample size equal to a power of $2$. Even an integrand with a QMC-friendly axis-parallel discrepancy is not generally integrated with an error of this rate. By using an importance sampling (Owen, 2013, Chapter 9) adjustment, we can arrange to obtain $n_{\ell}=2^{m_{\ell}}$ points from $P_{\ell}$ while also having $n=\sum_{\ell=1}^{L}n_{\ell}$ be a power of $2$. This can then give the $O(n_{\ell}^{-3+\epsilon})$ rate within every stratum and hence overall.
This paper is organized as follows. Section 2 introduces our notation, gives some background on RQMC, and proves some results about within-stratum discrepancies. Section 3 describes how to allocate samples to strata. We see that in RQMC sampling, it is advantageous to use more nearly equal allocations than those given by the stratum sizes. Section 4 shows how to partition a sample size $n=2^{m}$ into $L$ within-stratum sample sizes $n_{\ell}=2^{m_{\ell}}$. Section 5 establishes some minimax variance results from using equal or nearly equal sample sizes for the strata. Section 6 shows a simple toy example with $L=8$ strata where RQMC with $n_{\ell}$ taken to be powers of two does best. Section 7 uses a mixture model version of the Saint-Venant flood example from Limbourg and De Rocquigny (2010) where non-nominal sampling proportions do best. Section 8 has some discussion and conclusions. The proof of one of our results is in Appendix A.
@@
Klebanov and Sullivan (2023) describe using transport maps from a $\mathcal{N}(0,I)$ reference distribution to sample a mixture like $P=\sum_{\ell=1}^{L}\alpha_{j}\mathcal{N}(\mu_{j},\Sigma_{j})$. They construct a continuous mapping over $t\in[0,1]$ with $\mathcal{N}(0,I)$ at $t=0$ reaching $P$ at $t=1$ and solve an ordinary differential equation along that path. For $\ell<L$, they take the first $n_{\ell}=\lfloor n\alpha_{\ell}\rfloor$ points of a quasi-Monte Carlo sequence and use those to sample from component $\ell$. Then $n_{L}=n-\sum_{\ell=1}^{L-1}n_{\ell}$. This is different from our strategy of using the first RQMC component to select the stratum which we prefer because it avoids giving two within-stratum samplers the same points. They do not require $n\gg L$. Some of our expressions only require $n>2L$, but as we use asymptotic variances within strata, our results are most relevant for large $n/L$. They consider more general location-scale families than the Gaussian distribution and they can even include some $\alpha_{\ell}<0$ so long as $P$ never has a negative probability density function.
@@ 2.1 Integration variables
We will use RQMC points $\bm{z}\in[0,1]^{s+1}$. We partition $\bm{z}$ into its first component $v$ that we use to select a stratum $\ell$ and the subsequent $s$ components $\bm{u}$ that we use to sample $\bm{x}\sim P_{\ell}$. That is, $\bm{z}=(v,u_{1},u_{2},\dots,u_{s})$. For a function $\bm{\ell}:[0,1]\to\{1,2,\dots,L\}$, we will sample the $i$’th point using sample stratum $\bm{\ell}(v_{i})$. The pre-image of each $\ell\in\{1,2,\dots,L\}$ under $\bm{\ell}$ will always be a sub-interval of $[0,1]$. We let $\beta_{\ell}$ be the length of that subinterval. We abbreviate $\bm{\ell}(v_{i})$ to $\ell(i)$ in some expressions. Having selected stratum $\ell(i)$, we draw $\bm{x}_{i}\in^{d}$ via $\bm{x}_{i}=\phi_{\ell(i)}(\bm{u}_{i})$.
@@
is a within-stratum sample mean. For the RQMC method we consider, we show below that $n_{\ell}\geqslant 1$ whenever $\beta_{\ell}>2/n$. We do not need $\Pr(n_{\ell}>0)=1$ for unbiasedness of $\hat{\mu}$, but without it the stratum sample means $\hat{\mu}_{\ell}$ are not always well defined.
At this point we can introduce a small generalization. Since we already have a dependence on $\ell$ in $h_{\ell}$ we could also replace $h_{\ell}(\bm{u})=g(\phi_{\ell}(\bm{u}))$ by $h_{\ell}(\bm{u})=g_{\ell}(\phi_{\ell}(\bm{u}))=g_{\ell}(\bm{x})$ to allow the function being applied to $\bm{x}$ to vary with $\ell$.
@@ 2.2 Sampling the strata
We will assume that the sampling values $v_{1},\dots,v_{n}$ that we use are ‘stratified’. By this we mean that every interval $I_{k}=[(k-1)/n,k/n)$ contains precisely one of $v_{1},\dots,v_{n}$ for $k=1,\dots,n$. The next proposition shows that if every $\beta_{\ell}>2/n$, then every $n_{\ell}>0$ when $v_{i}$ are stratified.
@@
Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$. Then $\beta=t/n+c$ for $t=s-r$ and $-1/n<c<1/n$. If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (6) holds trivially.
@@ 2.4 Digital nets
For integers $m\geqslant t\geqslant 0$ the $n=b^{m}$ points $\bm{z}_{1},\dots,\bm{z}_{n}\in[0,1)^{d}$ are a $(t,m,d)$-net in base $b$ if $H(\bm{c},\bm{k})$ contains precisely $nb^{-|\bm{k}|}=b^{m-|\bm{k}|}$ of them, whenever $m-|\bm{k}|\geqslant t$. Smaller $t$ are better. We always have $t\leqslant m$ because $t=m$ just means that all the points are in $[0,1)^{d}$.
@@
While convergence $O(n^{-\rho})$ for $\rho\in\{1,2,3\}$ are widely studied, these come from error bounds and the actual convergence rate need not be near an integer. Liu (2025) shows examples where certain singular behavior of the integrand can give any rate between $1$ and $2$. In applications, the precise rate of convergence might not be known before computation begins.
@@ 2.5 Discrepancy within each stratum
Here we consider how uniformly distributed the points in each stratum are. The first case we consider has $\beta_{\ell}$ decreasing with $\ell$ and equal to negative powers of $2$. We show how to find such $\beta_{\ell}$ in Section 4. Then the sample points within any given stratum are scrambled nets, and the error can be as good as $O(n^{-3+\epsilon})$ for regular integrands. After that, we consider arbitrary $\beta_{\ell}>2/n$. Then the points in a given stratum still have small discrepancy but they are not generally scrambled digital nets. In that case the variance rate in stratum $\ell$ is $O(n_{\ell}^{-2+\epsilon})$ for within-stratum integrands that are in BVHK.
@@
If $m-\kappa_{\ell}-|\bm{k}|\geqslant t$, then by the digital net property of $\bm{z}_{1},\dots,\bm{z}_{n}$, this number is $2^{m-\kappa_{\ell}-|\bm{k}|}$. That is, if $m_{\ell}\geqslant|\bm{k}|+t$, then the number of such points is $2^{m_{\ell}-|\bm{k}|}$. ∎
@@ 3 Choosing $\beta_{\ell}$ for given $\alpha_{\ell}$
Values of $\rho$ equal to $1$, $2$ and $3$ correspond to upper bounds on the convergence rates for MC, QMC and RQMC variances, assuming $f\in L^{2}$, $V_{{\mathrm{HK}}}(f)<\infty$ and the aforementioned regularity condition on $f$, respectively, ignoring logarithmic factors in the latter two cases. It often happens that the empirical MSE of an RQMC rule follows an apparent rate of $n^{-\rho}$ for a value of $\rho$ that is not close to an integer. Unbounded integrands are not in BVHK, and yet they will have an MSE that is $o(n^{-1})$. Those often appear to have $21<\rho<$ 2.
@@
Either way, we can assume that $n_{\ell}=n\beta_{\ell}$. Then with $\omega_{\ell}=\alpha_{\ell}/\beta_{\ell}$, we can rewrite (5) as
@@
The importance sampling ratio $\omega_{\ell}$ has canceled from the expression for $\hat{\mu}$, leaving a simply weighted sum of within-stratum means.
@@ 3.1 Optimal sample sizes
RQMC points have a space filling property. By construction, the values of $\bm{u}_{i}\in[0,1]^{s}$ sampled for $\bm{\ell}(v_{i})=\ell$ tend to fill in the gaps left behind by $\bm{u}_{i}$ for $\bm{\ell}(v_{i})\neq\ell$. We might then anticipate primarily non-positive correlations among the $\hat{\mu}_{\ell}$. This motivates a different criterion
@@
In both criteria (12) and (14), higher convergence rates $\rho$ tend to equalize the $n_{\ell}$. We can think of $n_{\ell}=n/L$ as arising in the limit as $\rho\to\infty$. We will also see equal $n_{\ell}$ in a minimax argument in Section 5.
We now have six principled ways to choose how $n_{\ell}$ should scale with $\alpha_{\ell}$, for ansatzes 00 and $1$, and three different hypotheses on $\rho$. They lead to the four different rates given in Table 1. The case for ansatz 00 and $\rho=1$ (uncorrelated estimates at the MC rate) is the only one with a proportional allocation. The other choices all allocate more than a proportional number of samples to the strata with smaller $\alpha_{\ell}$. This is also true for the plain MC rate ($\rho=1$) under the conservative ansatz 1.
When $\alpha_{1}=\cdots=\alpha_{L}=1/L$, then equations (12) and (14) do not distinguish between different values of $\rho$. After raising $\alpha_{\ell}$ to any power they are still equal and then a near equal integer allocation is appropriate.
When we do have information about the relative sizes of $\tau_{\ell}$, then we can replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}$ in the formulas in Table 1. In that case, the proportionality for $C_{0}$ and $\rho=1$ matches the Neyman allocation from survey sampling, in Neyman (1934). When the cost to sample in stratum $\ell$ is proportional to $c_{\ell}>0$ and we constrain the total cost $\sum_{\ell=1}^{L}c_{\ell}n_{\ell}$, then the Lagrange multiplier argument leads us to replace $\alpha_{\ell}$ by $\alpha_{\ell}\tau_{\ell}^{1/2}/c_{\ell}$ in Table 1.
@@
If $\rho$ varies from stratum to stratum, taking the value $\rho_{\ell}$ in stratum $\ell$, then the Lagrange multiplier argument leads to $n_{\ell}\propto\alpha_{\ell}^{2/(\rho_{\ell}+1)}$. Strata with higher $\rho_{\ell}$ get larger $n_{\ell}$ and $\beta_{\ell}=n_{\ell}/n$ does not depend on $n$.